# KROMA: Ontology Matching with Knowledge Retrieval and Large Language Models

### v1

@@ 19                   $P\leftarrow(P\setminus\{C\})\cup\{C_{1},C_{2}\}$;
The offline refinement algorithm is outlined as Algorithm 1. (1) It starts by initializing ${\mathcal{G}}_{O}$ (lines 1-6) as the union of $O_{s}$ and $O_{t}$, followed by computing the node ranks. At each rank, it initializes a “bucket” $B_{i}$ (as a single node set) that simply include all the concept nodes at rank $i$ (lines 7-8), and initialize a partition ${\mathcal{P}}$ with all the buckets (line 9). It then follows a “bottom-up” process to refine the buckets iteratively, by checking if two concepts $c$ and $c^{\prime}$ in a same bucket $B_{i}$ are concept similar (as asserted by $\mathsf{LLM}$ and embedding similarity), and have all the neighbors that satisfy the requirement of ontological bisimilar relation by definition (lines 10-13). If not, a procedure $\mathsf{collapse}$ is invoked, to (1) split the bucket $B_{i}$ into three fragments: $B^{1}_{i}$ = $B_{i}\setminus\{c,c^{\prime}\}$, $B^{2}_{i}$ = $\{c\}$, and $B^{3}_{i}$=$\{c^{\prime}\}$, followed by a “merge” check to test if $c$ and $c^{\prime}$ can be merged to $B^{1}_{i}$; and (2) propagate this change to further “split-merge” operators to affected buckets at higher ranks (lines 14-17). This process continues until no change can be made.
@@
Time Cost. The initialization of concept graph ${\mathcal{G}}_{O}$ is in $O(|O_{s}|+|O_{t}|)$. Here $|O_{s}|$=$|V_{s}|+|E_{s}|$; and $|O_{t}|$ is defined similarly. The iterative collapse (lines 10-17) takes in $O(|O_{s}|+|O_{t}|)$ time as the number of buckets (resp. edges) is at most $|C_{s}|+|C_{T}|$ (resp. $|E_{s}|$+$|E_{t}|$). The overall cost is in $O(|O_{s}|+|O_{t}|)$.
@@
The algorithm (with pseudoscope reported in [1]) first updates the buckets in ${\mathcal{G}}_{O}$ by incorporating the nodes from $\Delta{\mathcal{G}}$ that are verified to be concept similar, as well as their ranks. It then incrementally update the buckets to maintain the bisimilar equivalence consistently via a “bottom-up” split-merge process as in Algorithm 1 (lines 6-15). Due to the unpredictability of the “unseen” concepts, the online refinement defers the processing of two “inconsistent” cases for experts’ validation: (1) When a concept $c$ is determined to be concept similar by function $\sim$ but not $\mathsf{LLMs}$ with high confidence; or (2) whenever for a new edge $(c,c^{\prime})\in\Delta{\mathcal{G}}$, $(c,[c_{1}])\in R_{\simeq}$, $(c^{\prime},[c_{2}])\in R_{\simeq}$, yet $r(c_{1})<r(c_{2})$ in ${\mathcal{G}}_{O}$. Both require domain experts’ feedback to resolve. These cases are cached into a query set ${\mathcal{Q}}$ to be further resolved in the validation phase (see Section 4). We cache these cases into a query set by using an auxiliary data structure (e.g., priority queue ranked by $\mathsf{LLM}$ confidence score) to manage their processing.
Analysis. The correctness of online refinement carries over from its offline counterpart, and that it correctly incrementalize the split-merge operator for each newly arrived edges. For time cost, for each batch, it takes a delay time to update ${\mathcal{G}}_{O}$ in $O(|{\mathcal{G}}_{O}|+|\Delta G|\log|\Delta G|+|\Delta G|\log|V_{O}|)$ time. This result verifies that online refinement is able to response faster than offline maintenance that recomputes the concept graph from scratch. We present the detailed analysis in [1].
@@ 6 Experimental Study
We investigated the following research questions. [RQ1]: How well $\mathsf{KROMA}$ improves state-of-the-art baselines with different $\mathsf{LLMs}$? [RQ2]: How can knowledge retrieval and ontology refinement enhance matching performance? and [RQ3]: What are the impact of alternative $\mathsf{LLM}$ reasoning strategies?
@@ 6.1 Experimental Setting
Benchmark Datasets. We selected five tracks from the OAEI campaign [41], covering various domain tracks. For each track, we selected two representative ontologies as a source ontology $O_{s}$ and a target ontology $O_{t}$. The selected tracks include Anatomy [16] (Mouse-Human), Bio-LLM [24] (NCIT-DOID), CommonKG (CKG) [19] (Nell-DBpedia, YAGO-Wikidata), BioDiv [29] (ENVO-SWEET), and MSE [26] (MI-MatOnto). To ensure a fair and comprehensive evaluation of KROMA, we adopted the standard benchmarks from the Ontology Alignment Evaluation Initiative (OAEI), enabling direct comparison with prior LLM-based methods. Despite the high cost of LLM inference, we tested KROMA across five diverse tracks to demonstrate its robustness across domains.
$\mathsf{LLMs}$ selection. To underscore $\mathsf{KROMA}$ ’s ability to achieve strong matching performance even with smaller or lower-performance LLMs, we selected models with relatively modest Chatbot Arena MMLU scores [9]: Gemma-2B (51.3%) and Llama-3.2-3B (63.4%), compared to the baseline systems Flan-T5-XXL (55.1%) and MPT- 7B (60.1%). We have selected a diverse set of $\mathsf{LLMs}$, ranging from ultra-lightweight to large-scale—to demonstrate $\mathsf{KROMA}$’s compatibility with models that can be deployed on modest hardware without sacrificing matching quality. Our core evaluations use DeepSeek-R1-Distill-Qwen-1.5B [51] (1.5B), GPT-4o-mini [36] (8B), and Llama-3.3 [31] (70B), each chosen in a variant smaller than those employed by prior OM-LLM baselines. To further benchmark our framework, we include Gemma-2B [11] (2B), Llama-3.2-3B [3] (3B), Mistral-7B [4] (7B), and Llama-2-7B [2] (7B) in our ablation studies. To run inference on the aforementioned models, we make use of TogetherAI and OpenAI APIs as off-the-shelf inference services on hosted, pretrained models, and we are not performing any fine-tuning or weight updates.
$\mathsf{LLMs}$ Determinism & Reproducibility. To enforce determinism and ensure reproducibility, all LLM calls are issued with a fixed temperature of 0.3, striking a balance between exploration and exploitation; and a constant random seed is used for sampling.
@@
Baselines.
Our evaluation considers four state-of-the-art $\mathsf{LLM}$-based ontology matching methods: LLM4OM [20], MILA [49], OLaLa [25], and LLMap [23]. For RQ2, we also developed KROMA-NR, which skip knowledge retrieval, and KROMA-NB, which disables the bisimilarity-based clustering (hence clusters are determined by concept similarity alone).
Naming Convention. Each configuration uses a pattern [Method][Optional Suffix][LLM Version], where the initials (K, M, O, L) specify the method (KROMA, MILA, OLaLa, LLM4OM). Suffixes “NKR” and “NR” denote “no knowledge retrieval” and “no ontology refinement”, respectively, and the trailing version number (e.g., 3.3, 2.0, 4mini) specifies the underlying LLM release. Some examples are: KL3.3 is the full $\mathsf{KROMA}$ using LLaMA-3.3; KNR3.3 is $\mathsf{KROMA}$ without Knowledge Retrieval using LLaMA-3.3; and KNB3.3 is KROMA without Ontology Refinement using LLaMA-3.3. The convention applies similarly to other $\mathsf{LLMs}$ (e.g., Gemma-2B, Mistral-7B, GPT-4o-mini) and $\mathsf{LLM}$-based ontology matching methods (MILA, OLaLa, LLM4OM).
@@ 6.2 Experimental Results
$\mathsf{KROMA}$ vs. Baselines: Overall Performance. Across all six datasets in Table 2 (abbreviated by their first capitalized letters), the full $\mathsf{KROMA}$ configuration (KL3.3) achieves the highest $F_{1}$ on every task, substantially outperforming competing baselines. For Mouse–Human, KL3.3 delivers 94.94 $F_{1}$, eclipsing MILA’s best (ML3.1) at 92.20, OLaLa (OL2.0) at 90.20, and LLM4OM (L4G3.5) at 89.11. On NCIT–DOID, KL3.3 reaches 98.63 versus 94.80 for MILA, 83.01 for LLM4OM, and 72.10 for Flan-T5-XXL. Similar gaps appear on Nell–DBpedia (97.08 vs. 96.00/94.26), YAGO–Wikidata (95.54 vs. 92.19/93.33), ENVO–SWEET (93.18 vs. 92.98/85.06), and MI–MatOnto (61.25 vs. 59.05/32.97). These verify the effectiveness of $\mathsf{KROMA}$ over representative benchmark datasets. More details (e.g., Precision and Recall) are reported in [1].
@@
Impact of Test sizes. We next report the impact of ontology sizes to the performance of $\mathsf{KROMA}$ in performance (detailed results reported in [1]). For each dataset, we varied test sizes from 200 (xsmall) to 1,000 (full) pairs. $\mathsf{KROMA}$’s performance is in general insensitive to the change of test sizes. For example, its $F_{1}$ stays within a 1.90 variance on NCIT–DOID and a 1.10-point range on YAGO–Wikidata. This verifies the robustness of $\mathsf{KROMA}$ in maintaining desirable performance for large-scale ontology matching tasks.
@@
Table 4: $\mathsf{KROMA}$ Performance with/without Ontology Refinement (Left); and with/without Knowledge Retrieval (Right).
@@
Impact of Ontology Refinement. Table 4 (Left) shows that by incorporating Ontology Refinement ($\mathsf{KROMA}$ vs. $\mathsf{KROMA}$-NR), $\mathsf{KROMA}$ improves $F_{1}$ score across 6 datasets: Mouse–Human (+3.69), NCIT–DOID (+3.62), Nell–DBpedia (+2.82), YAGO–Wikidata (+3.56), ENVO–SWEET (+1.06), MI–MatOnto (+1.30). By pruning false candidate pairs via offline and online refinement strategies on the concept graph, ontology refinement retains true semantic connections, by filtering out noise while preserving true alignments.
@@
Exp-3: Alternative $\mathsf{LLM}$ Reasoning Strategies. We evaluate $\mathsf{KROMA}$’s performance with two alternative $\mathsf{LLM}$ reasoning strategies: “Debating” and “Deep reasoning” to understand their impact to ontology matching.
@@
Dataset Model P R F1 NCIT–DOID L3.3Short 97.59 99.69 98.63 L3.3Long 97.66 96.31 96.98 Nell–DBpedia L3.3Short 94.32 100.00 97.08 L3.3Long 94.67 95.48 95.07 YAGO–Wikidata L3.3Short 91.80 99.60 95.54 L3.3Long 92.31 94.80 93.53
@@ 7 Conclusion
Supplemental Material Statement. Our code and experimental data has been made available at https://anonymous.4open.science/r/kroma/, including an extended version of the paper [1].

### v2

@@ 19       $P\leftarrow(P\setminus\{C\})\cup\{C_{1},C_{2}\}$;
The offline refinement algorithm is outlined as Algorithm 1. (1) It starts by initializing ${\mathcal{G}}_{O}$ (lines 1-6) as the union of $O_{s}$ and $O_{t}$, followed by computing the node ranks. At each rank, it initializes a “bucket” $B_{i}$ (as a single node set) that simply include all the concept nodes at rank $i$ (lines 7-8), and initialize a partition $\mathcal{P}$} with all the buckets (line 9). It then follows a “bottom-up” process to refine the buckets iteratively, by checking if two concepts $c$ and $c^{\prime}$ in a same bucket $B_{i}$ are concept similar (as asserted by $\mathsf{LLM}$ and embedding similarity), and have all the neighbors that satisfy the requirement of ontological bisimilar relation by definition (lines 10-13). If not, a procedure $\mathsf{collapse}$ is invoked, to (1) split the bucket $B_{i}$ into three fragments: $B^{1}_{i}$ = $B_{i}\setminus\{c,c^{\prime}\}$, $B^{2}_{i}$ = $\{c\}$, and $B^{3}_{i}$$\{c^{\prime}\}$, followed by a “merge” check to test if $c$ and $c^{\prime}$ can be merged to $B^{1}_{i}$; and (2) propagate this change to further “split-merge” operators to affected buckets at higher ranks (lines 14-17). This process continues until no change can be made.
@@
Time Cost. The initialization of concept graph ${\mathcal{G}}_{O}$ is in $O(|O_{s}|+|O_{t}|)$. Here $|O_{s}|$$|V_{s}|+|E_{s}|$; and $|O_{t}|$ is defined similarly. The iterative collapse (lines 10-17) takes in $O(|O_{s}|+|O_{t}|)$ time as the number of buckets (resp. edges) is at most $|C_{s}|+|C_{T}|$ (resp. $|E_{s}|$$|E_{t}|$). The overall cost is in $O(|O_{s}|+|O_{t}|)$.
@@
The algorithm (with pseudoscope reported in [1]) first updates the buckets in ${\mathcal{G}}_{O}$ by incorporating the nodes from $\Delta{\mathcal{G}}$ that are verified to be concept similar, as well as their ranks. It then incrementally update the buckets to maintain the bisimilar equivalence consistently via a “bottom-up” split-merge process as in Algorithm 1 (lines 6-15). Due to the unpredictability of the “unseen” concepts, the online refinement defers the processing of two “inconsistent” cases for experts’ validation: (1) When a concept $c$ is determined to be concept similar by function $\sim$ but not $\mathsf{LLMs}$ with high confidence; or (2) whenever for a new edge $(c,c^{\prime})\in\Delta{\mathcal{G}}$, $(c,[c_{1}])\in R_{\simeq}$, $(c^{\prime},[c_{2}])\in R_{\simeq}$, yet $r(c_{1})<r(c_{2})$ in ${\mathcal{G}}_{O}$. Both require domain experts’ feedback to resolve. These cases are cached into a query set $\mathcal{Q}$} to be further resolved in the validation phase (see Section 4). We cache these cases into a query set by using an auxiliary data structure (e.g., using a priority queue ranked by $\mathsf{LLM}$ confidence scores) to effectively manage their processing.
Analysis. The correctness of online refinement carries over from its offline counterpart, and that it correctly incrementalize the split-merge operator for each newly arrived edges. For time cost, for each batch, it takes a delay time to update ${\mathcal{G}}_{O}$ in $O(|{\mathcal{G}}_{O}|+|\Delta G|\log|\Delta G|+|\Delta G|\log|V_{O}|)$ time. This result verifies that online refinement is able to response faster than offline maintenance that recomputes the concept graph from scratch. We present detailed analysis in [1].
@@ 6 Experimental Study
We investigated the following research questions. [RQ1]: How well $\mathsf{KROMA}$ improves state-of-the-art baselines with different $\mathsf{LLMs}$? [RQ2]: How can knowledge retrieval and ontology refinement enhance matching performance? and [RQ3]: What are the impact of using alternative $\mathsf{LLM}$ reasoning strategies?
@@ 6.1 Experimental Setting
Benchmark Datasets. We selected five tracks from the OAEI campaign [32], covering various domain tracks. For each track, we selected two representative ontologies as a source ontology $O_{s}$ and a target ontology $O_{t}$. The selected tracks include Anatomy [16] (Mouse-Human), Bio-LLM [23] (NCIT-DOID), CommonKG (CKG) [19] (Nell-DBpedia, YAGO-Wikidata), BioDiv [28] (ENVO-SWEET), and MSE [25] (MI-MatOnto). To ensure a fair and comprehensive evaluation of KROMA, we adopted the standard benchmarks from the Ontology Alignment Evaluation Initiative (OAEI), enabling direct comparison with prior LLM-based methods. Despite the high cost of LLM inference, we tested KROMA across five diverse tracks to demonstrate its robustness and effectiveness across domains.
$\mathsf{LLMs}$ selection. To underscore $\mathsf{KROMA}$ ’s ability to achieve strong matching performance even with smaller or lower-performance LLMs, we selected models with relatively modest Chatbot Arena MMLU scores [9]: Gemma-2B (51.3%) and Llama-3.2-3B (63.4%), compared to the baseline systems Flan-T5-XXL (55.1%) and MPT- 7B (60.1%). We have selected a diverse set of $\mathsf{LLMs}$, ranging from ultra-lightweight to large-scale—to demonstrate $\mathsf{KROMA}$’s compatibility with models that can be deployed on modest hardware without sacrificing matching quality. Our core evaluations use DeepSeek-R1-Distill-Qwen-1.5B [48] (1.5B), GPT-4o-mini [34] (8B), and Llama-3.3 [21] (70B), each chosen in a variant smaller than those employed by prior OM-LLM baselines. To further benchmark our framework, we include Gemma-2B [11] (2B), Llama-3.2-3B [47] (3B), Mistral-7B [3] (7B), and Llama-2-7B [2] (7B) in our ablation studies. To run inference on the aforementioned models, we used TogetherAI and OpenAI APIs, treating them as off-the-shelf inference services on hosted, pretrained models. We are not performing any fine-tuning or weight updates to the models.
@@
Baselines. Our evaluation considers four state-of-the-art $\mathsf{LLM}$-based ontology matching methods: LLM4OM [20], MILA [45], OLaLa [24], and LLMap [23]. For RQ2, we also developed KROMA-NR, which skips the knowledge retrieval process, and KROMA-NB, which disables the use of bisimilarity-based clustering, allowing only the use of concept similarity to determine concept clusters.
Naming Convention. Each configuration uses a pattern [Method][Optional Suffix][LLM Version], where the initials (K, M, O, L) specify the method (KROMA, MILA, OLaLa, LLM4OM). Suffixes “NKR” and “NR” denote “no knowledge retrieval” and “no ontology refinement”, respectively, and the trailing version number (e.g., 3.3, 2.0, 4mini) specifies the underlying LLM release.
@@ 6.2 Experimental Results
$\mathsf{KROMA}$ vs. Baselines: Overall Performance. Across all six datasets in Table 2 (abbreviated by their first capitalized letters), the full $\mathsf{KROMA}$ configuration (KL3.3) achieves the highest $F_{1}$ on every task, substantially outperforming competing baselines. For Mouse–Human, KL3.3 delivers 94.94 $F_{1}$, eclipsing MILA’s best (ML3.1) at 92.20, OLaLa (OL2.0) at 90.20, and LLM4OM (L4G3.5) at 89.11. On NCIT–DOID, KL3.3 reaches 98.63 versus 94.80 for MILA, 83.01 for LLM4OM, and 72.10 for Flan-T5-XXL. Similar gaps appear on Nell–DBpedia (97.08 vs. 96.00/94.26), YAGO–Wikidata (95.54 vs. 92.19/93.33), ENVO–SWEET (93.18 vs. 92.98/85.06), and MI–MatOnto (61.25 vs. 59.05/32.97). These cumulative results verify the effectiveness of $\mathsf{KROMA}$ over representative benchmark datasets. More details (e.g., Precision and Recall) are reported in [1].
@@
Impact of Test sizes. We next report the impact of ontology sizes to the performance of $\mathsf{KROMA}$ in performance (detailed results reported in [1]). For each dataset, we varied test sizes from 200 (xsmall) to 1,000 (full) pairs. $\mathsf{KROMA}$’s performance is in general insensitive to the change of test sizes. For example, its $F_{1}$ stays within a 1.90 variance on NCIT–DOID and a 1.10-point range on YAGO–Wikidata. This verifies the robustness and effectiveness of $\mathsf{KROMA}$ in maintaining desirable performance for large-scale ontology matching tasks.
@@
Table 4: $\mathsf{KROMA}$ performance under different scenarios: with/without Ontology Refinement (Left); and with/without Knowledge Retrieval (Right).
@@
Impact of Ontology Refinement. Table 4 (Left) shows that by incorporating Ontology Refinement ($\mathsf{KROMA}$ vs. $\mathsf{KROMA}$-NR), $\mathsf{KROMA}$ improves $F_{1}$ score across 6 datasets: Mouse–Human (+3.69), NCIT–DOID (+3.62), Nell–DBpedia (+2.82), YAGO–Wikidata (+3.56), ENVO–SWEET (+1.06), MI–MatOnto (+1.30). By pruning false candidate pairs using offline and online strategies on the concept graph, $\mathsf{KROMA}$ is able to retain true semantical connections between concepts.
@@
Exp-3: Alternative $\mathsf{LLM}$ Reasoning Strategies. We evaluated $\mathsf{KROMA}$’s performance with two alternative $\mathsf{LLM}$ reasoning strategies: “Debating” and “Deep reasoning” to further understand their impact to ontology matching.
@@
Dataset Model P R F1
NCIT–DOID L3.3Short 97.59 99.69 98.63
L3.3Long 97.66 96.31 96.98
Nell–DBpedia L3.3Short 94.32 100.00 97.08
L3.3Long 94.67 95.48 95.07
YAGO–Wikidata L3.3Short 91.80 99.60 95.54
L3.3Long 92.31 94.80 93.53
@@
Can “Active Sampling” help? In the semi-supervised experiments on the Bio-LLM NCIT-DOID benchmark, KROMA paired with Llama-3.3-70B matches or exceeds the performance of MILA + Llama-3.1-70B. Under an active-learning regime with few-shot demonstrations, our method scores 1.90 F1 points higher than MILA and surpasses the previous state of the art by over 6 points. As Table 6 shows, combining bisimilarity-guided refinement with selective querying delivers a substantial boost in $\mathsf{KROMA}$’s alignment performance.
Table 6: Active Learning on NCIT-DOID.
Dataset Model $P$ $R$ $F_{1}$
NCIT-DOID KROMA + Llama-3.1-70B 97.85 98.15 98.00
MILA + Llama-3.1-70B 96.70 92.80 94.67
@@ 7 Conclusion
Supplemental Material Statement. Our code and experimental dataset has been made available at https://github.com/lamng3/kroma, including an extended version of the paper [1], providing further details for $\mathsf{KROMA}$.
Acknowledgements. This material is based upon research in the Materials Data Science for Stockpile Stewardship Center of Excellence (MDS3-COE), and supported by the Department of Energy’s National Nuclear Security Administration under Award Number(s) DE-NA0004104. This material is based upon work supported by the U.S. National Science Foundation, Office of Advanced Cyberinfrastructure (OAC), Major Research Instrumentation, under Award Number 2117439. This work made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at CWRU.